OpenStudy (anonymous):

Evaluate the integral Integral of cost/(1+(sint)^2)^1/2 from 0 to pi over 2

OpenStudy (amistre64):

seeing that cos is the derivative of sin; youd think there was a chain rule involved

OpenStudy (amistre64):

\[\int\frac{cos(t)}{(1+sin^2(t))^{1/2}}dt \] \[\int cos(t)(1+sin^2(t))^{-1/2}dt \] what happens when we try to derive something close to it? \[\frac{d}{dx}\ (1+sin^2(t))^{1/2}\] \[\frac{(1+sin^2(t))'}{2(1+sin^2(t))^{1/2}}\] \[\frac{2\ sin(t)\ cos(t)}{2(1+sin^2(t))^{1/2}}\] \[\frac{sin(t)\ cos(t)}{(1+sin^2(t))^{1/2}}\] which is close, but is it helpful?

OpenStudy (jamesj):

Substitute u = sin t And you'll get the integral of 1/sqrt(1+u^2) Now make a second substitution, u = sinh t (that's hyperbolic sin) and you'll get the integral of 1 dt

OpenStudy (amistre64):

my other thought is to use an ln(...) config

OpenStudy (amistre64):

doh!! ... hyperbolics

OpenStudy (anonymous):

Thank you, this has all been very helpful.